OpenStudy (anonymous):

Write a system that describe the situation: Alex purchased 2 hotdogs for x dollars each and 4 popcorn balls for y dollars each for a total cost of $14.00. Marion purchased 4 hotdogs for x dollars each and 7 popcorn balls for y dollars each for a total $26.00.

OpenStudy (anonymous):

2x+4y=14 and 4x+7y=26

OpenStudy (anonymous):

Could you explain how you got that please?

OpenStudy (anonymous):

id alex purchased 2 hotdogs for x dollars since we don't know the value of x or y we set it as 2x+4y=14 because the cost time the amount of popcorm and hotdog for the first one is 14. 2)same thing

OpenStudy (anonymous):

Thank you!